OpenStudy (anonymous):

write 4716 in scientific notation

OpenStudy (anonymous):

@jdoe0001 last 2 questions I promise

OpenStudy (jdoe0001):

http://www.mathsisfun.com/numbers/scientific-notation.html scientific notation means representation with a base of 10 and I think it has to have only 1 number left of the decimal point so you have 4 numbers so we'll move 3 of them to the right of the decimal point that means, we'll use 3 also in the base 10 or \(\large \bf 4\underline{716}\implies 4.716\times 10^3\)

OpenStudy (jdoe0001):

\(\bf 4.716\times 10^3\implies 4.716\times 1000\implies 4716\) thus

OpenStudy (anonymous):

so the scientific notation of 4716 = 4.716 X 10^3?

OpenStudy (jdoe0001):

yeap